OpenStudy (anonymous):

Which ratio is equivalent to 9/27 ? Option A: 6/9 Option B: 1/2 Option C: 1/3 Option D: 3/12

OpenStudy (calculusxy):

C

OpenStudy (anonymous):

C! :) 9*3=27

OpenStudy (anonymous):

Thank you guys so much! :D

OpenStudy (calculusxy):

9/9=1 27/9=3 1/3